1/y-1+5/12=-2/3y-3
Answer by fractalier(6550)

We multiply all by the lowest common denominator to clear fractions...in this case 12(y - 1)...so we get
1/y-1 + 5/12 = -2/3y-3
12(y - 1)[1/y-1 + 5/12 = -2/3y-3]
12 + 5(y-1) = -8
12 + 5y - 5 = -8
5y = -15
y = -3